I have a quiz tomorrow and this is a problem on my homework, I could really use some help!
A chemist has a 60% acid solution and a 20% acid solution. How many millimeters of each should be used to make 140ml of a 35% acid solution? Let x represent the number of ml of the 60% acid solution.
Thanks so much!!

Found 2 solutions by josgarithmetic, stanbon:
Answer by josgarithmetic(39617)
Answer by stanbon(75887)

A chemist has a 60% acid solution and a 20% acid solution. How many millimeters of each should be used to make 140ml of a 35% acid solution? Let x represent the number of ml of the 60% acid solution.
-----
Using two equations::
Quantity:: s + t = 140 ml
Acid Eq:::0.60s + 0.20t = 0.35*140
-------------------------------------------
Modify for elimination:
60s + 60t = 60*140
60s + 20t = 35*140
---------------------
Subtract and solve for "t"::
40t = 25*140
-----
t = 87.5 ml (amt. of 20% solution needed
---
s = 140-t = 52.5 ml (amt. of 60% solution needed)
----------------------------
OR
Using one equation:
acid + acid = acid
0.60x + 0.20(140-x) = 0.35*140
Multiply thru by 100 to get:
60x + 20*140 - 20x = 35*140
40x = 15*140
x = 52.5 ml (amt. of 60% solution needed)
140-x = 87.5 ml (amt. of 20% solution needed)
-----------
Cheers,
Stan H.
------------